OpenStudy (anonymous):

The series \[\large \frac{2x}{x+3}+(\frac{2x}{x+3})^{2}+(\frac{2x}{x+3})^{3}+...........\infty \] will have a non zero definite sum then find the complete x.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

ans. -1<x<3 (x not = to 0) (solution needed)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\large{1+n+n^2+n^3+\dots +\infty=\frac{1}{1-n}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then..

OpenStudy (experimentx):

let n = 2x/(x+3) also note that this only holds for case 2x/(x+3) < 1

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

using this 2x/(x+3) < 1 find the interval of convergence.

OpenStudy (anonymous):

(-3,3)

OpenStudy (anonymous):

@experimentX but ans.-1<x<3

OpenStudy (anonymous):

\[\left| \frac{2x}{x+3} \right|<1 \Rightarrow \left| \frac{x}{x+3} \right|<\frac{1}{2} \Rightarrow \left| 1-\frac{3}{x+3} \right|<\frac{1}{2} \Rightarrow -\frac{1}{2}< 1-\frac{3}{x+3}<\frac{1}{2}\\-\frac{3}{2}< -\frac{3}{x+3}<-\frac{1}{2}\\\frac{1}{2}< \frac{3}{x+3}<\frac{3}{2}\\\frac{1}{6}< \frac{1}{x+3}<\frac{1}{2}\\ 2<x+3<6 \\ -1<x<3 \]